Please help me solve this word problem.
Solve with using an algebraic equation an unknown.
The perimeter of a rectangle shed is 43 feet. It's length is 5 feet more than twice its width. Find the length and width of the shed.

Answer by checkley79(3341)

P=2L+2W
L=2W+5
43=2(2W+5)+2W
43=4W+10+2W
43=6W+10
6W=43-10
6W=33
W=33/6
W=5.5 ANS, FOR THE WIDTH.
L=2*5.5+5
L=11+5=16 ANS. FOR THE LENGTH.
PROOF:
43=2*16+2*5.5
43=32+11
43=43